OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (turingtest):

x=#two-steps y=#one-steps y+x=1070 3y-2x=110 multiply the first eqn. by 2 and add it to the second equation: 5y=2250 y=450

OpenStudy (anonymous):

\[3x_1-2x_2=110. solve for x\]

OpenStudy (turingtest):

@alexmiles you cant solve for x when you have 2 variables and only one eqn.

OpenStudy (anonymous):

* in combination with the other equation, solve for \[x_2\]